OpenStudy (anonymous):

Use the elimination method to solve the system of equations: a+7b=7 a+2b= -3

OpenStudy (zehanz):

Subtract the second equation from the first to eliminate a, then calculate b. At last because you have b, you can also find a!

OpenStudy (anonymous):

I am confused math is not my best subject

OpenStudy (zehanz):

Do this: (i) a+7b=7 (ii) a+2b=-3 -------- - (iii) 5b=10 Explanation: a-a eliminates a Also calculate 7b-2b=5b Finally 7-(-3)=7+3=10 So now you have 5b=10. Calculate b, then substitute that value in eq. (i) or (ii)

OpenStudy (anonymous):

would it look something like this \[a + 7b= 7\]\[a + 2b = -3\]

OpenStudy (anonymous):

I hope I dont sound stupid here but when calculating for b would I reverse the numbers so it would like b+7a=7 and then b+2a=-3

OpenStudy (zehanz):

No, once you've found what b is (solve 5b=10), use that value to replace b in a+7b=7.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

a+7(10)=7